a club is compose of 10members wish to elect officers from president to treasurer .how many sets of officer is possible?
Answer by nerdybill(7384)

a club is compose of 10members wish to elect officers from president to treasurer .how many sets of officer is possible?
.
Since order is important, we are talking about a "combination".
And, it depends on how many officers there are.
For example, if the officers are:
president
vice president
secretary
treasurer
.
Then:
10C4
= 10!/4!(10-4)!
= 10!/4!6!
= (7*8*9*10)/(1*2*3*4)
= (7*2*9*10)/(1*2*3)
= (7*9*10)/(3)
= 7*3*10
= 21*10
= 210